1 The Wildcard Module

Wildcards, also called “Shell Filename Patterns”, describe sets of file names. The wildcard package exports the following symbol:

Function match pattern string [:start] [:end] [:case-insensitive] This function returns a non-nil value if the string matches the pattern.

2 Wildcard Syntax

* Matches any zero or more characters. ? Matches any one character.

[string] Matches exactly one character that is a member of the string string. This is called a character class. As a shorthand, string may contain ranges, which consist of two characters with a dash between them. For example, the class ‘[a-z0-9_]’ matches a lowercase letter, a number, or an underscore. You can negate a class by placing a ‘!’ or ‘^’ immediately after the opening bracket. Thus, ‘[^A-Z@]’ matches any character except an uppercase letter or an at sign.

\ Removes the special meaning of the character that follows it. This works even in character classes.

Slash characters have no special significance in the wildcard matching, unlike in the shell, in which wildcards do not match them. Therefore, a pattern ‘foo*bar’ can match a file name ‘foo3/bar’, and a pattern ‘./sr*sc’ can match a file name ‘./src/misc’.